### v5

# A Unit Enhancement and Guidance Framework for Audio-Driven Avatar Video Generation

Abstract.
Audio-driven human animation technology is widely used in human-computer interaction, and the emergence of diffusion models has further advanced its development. Currently, most methods rely on multi-stage generation and intermediate representations, resulting in long inference time and issues with generation quality in specific foreground regions and audio-motion consistency. These shortcomings are primarily due to the lack of localized fine-grained supervised guidance. To address above challenges, we propose Parts-aware Audio-driven Human Animation, PAHA, a unit enhancement and guidance framework for audio-driven upper-body animation. We introduce two key methods: Parts-Aware Re-weighting (PAR) and Parts Consistency Enhancement (PCE). PAR dynamically adjusts regional training loss weights based on pose confidence scores, effectively improving visual quality. PCE constructs and trains diffusion-based regional audio-visual classifiers to improve the consistency of motion and co-speech audio. Afterwards, we design two novel inference guidance methods for the foregoing classifiers, Sequential Guidance (SG) and Differential Guidance (DG), to balance efficiency and quality respectively. Additionally, we build CNAS, the first public Chinese News Anchor Speech dataset, to advance research and validation in this field. Extensive experimental results and user studies demonstrate that PAHA significantly outperforms existing methods in audio-motion alignment and video-related evaluations. The codes and CNAS dataset will be released upon acceptance.
Human Animation, Video Generation, Parts-Aware, Audio-Driven
Figure 1. (a) S2G (He et al., 2024), the state-of-the-art for co-speech gesture generation, suffers from localized poor quality (e.g., hands, face) and audio-motion misalignment. (b) Qualitative ablation study. Parts-Aware Re-weighting (PAR) improves local generation quality of characters, while Parts Consistency Enhancement (PCE) enhances alignment between motion and co-speech audio. Our comprehensive method generates high-quality and consistent videos. (c) Comparison of baselines and our PAHA in terms of video quality (FVD, lower is better), diversity (Div., higher is better), and inference efficiency (TC(Time Cost), lower is better). Our method (PAHA-DG and PAHA-SG) achieves superior FVD and Div. performance while maintaining lower TC.
1. Introduction
Human-centered content generation has been a focal point in computer vision research. Human animation technology aims to synthesize speaking character videos from a single static reference character image and corresponding speech audio. This technology holds significant value across various fields, including video games, virtual reality, film and television production, social media, digital marketing, online education, human-computer interaction, and virtual assistants (Nazarieh et al., 2024).
Recent advances in GAN and diffusion models (Corona et al., 2024; Wei et al., 2024; Tian et al., 2025; Xu et al., 2024a; Song et al., 2022; Zhu et al., 2025; Zhang et al., 2024b; Yao et al., 2024; Liu et al., 2024) have enhanced high-resolution, high-quality character generation, particularly for maintaining long-term identity consistency. For instance, Live Portrait (Guo et al., 2024) uses GANs for portrait animation with stitching and redirection controls, while diffusion-based methods like VASA-1 (Xu et al., 2024a), EMO (Tian et al., 2025), and Hallo (Xu et al., 2024b) enable end-to-end human animation. However, these methods often struggle with poor generation quality in specific regions (e.g., lips, eyes) and focus primarily on audio-driven talking faces, neglecting gestures and body motions, which restricts their applicability. Actually, co-speech gesture generation (He et al., 2024; Liu et al., 2022a; Qian et al., 2021), treated as a separate task, highlights gestures’ role in enhancing speech clarity, persuasion, and credibility. For example, 3D-GestureGAN (Mahapatra et al., 2025) produces realistic talking videos with natural gestures using UV texture optimization and conditional GANs, while S2G (He et al., 2024) combines thin-plate-splines (TPS) transformations with a transformer-based diffusion model to generate long-duration, high-quality co-speech gesture videos. Despite progress, these methods face challenges in achieving speech-motion synchronization due to multi-stage training and inference, leading to higher computational costs and error accumulation. The reliance on intermediate representations, such as 2D/3D landmarks, TPS, 3D mesh, and optical flow, complicates processes further. These limitations primarily stem from the lack of localized fine-grained supervised learning or guidance.
Motivated by the divide-and-conquer strategy, we propose Parts-Aware Audio-Driven Human Animation (PAHA). Our method independently optimizes co-speech face and gesture generation in distinct spatial areas within a single video diffusion model, effectively addressing the limitations of existing methods regarding local poor quality and audio-video misalignment (Fig.1). Parts-Aware Re-weighting (PAR) (Sec.4.1.2) assigns dynamic loss weights to key areas (e.g., hands, face, body) in the training video frames. Amplifying regional loss enhances the model’s focus and learning in these areas, improving generation quality. Parts Consistency Enhancement (PCE) (Sec.4.1.3) trains diffusion-based regional classifiers to distinguish audio-visual differences between generated and real samples, learning the temporal correlation between spectral energy variations and localized character motion. During inference, classifiers provide gradient alignment signals to guide the model’s focus on specific regions, enhancing audio-video alignment. Subsequently, we design two innovative inference guidance methods for the aforementioned classifiers: Sequential Guidance (SG) (Sec.4.2.1) for efficiency and Differential Guidance (DG) (Sec.4.2.2) for quality.
Additionally, given the lack of open-source Chinese co-speech gesture datasets, we constructed the Chinese News Anchor Speech Dataset (CNAS). This dataset comprises videos featuring dynamic gestures, which highlights the complexities of human communication. Multilingual datasets are crucial for evaluating methods comprehensively. We validate its effectiveness against our method and baselines.
Our contributions are summarized as follows:
  • We propose PAHA, an end-to-end framework for generating audio-driven upper-body human animation. Parts-Aware Re-weighting (PAR) method dynamically adjusts loss based on keypoint confidence scores, improving animation quality.
  • We introduce Parts Consistency Enhancement (PCE), which uses self-distillation to train diffusion-based regional classifiers, enabling them to learn the temporal correlation between character motion and audio spectrum.
  • During inference, we apply classifier-based consistency guidance to align regional motion with audio, offering two approaches: Sequential Guidance (SG) for efficiency and Differential Guidance (DG) for quality.
  • We create the Chinese News Anchor Speech Dataset (CNAS) to address the lack of Chinese co-speech datasets and validate multiple methods on it.
  • Extensive experiments show that our framework produces high-quality, lifelike animations aligned with audio and outperforms previous state-of-the-art methods in quantitative and qualitative evaluations.
2. Related Work
2.1. Audio-Driven Human Animation
Audio-driven human animation methods can be divided into two categories: talking head generation and co-speech gesture generation. Talking head generation focuses on head motion and facial expression quality. These methods commonly use 3D meshes, 2D/3D landmarks, NeRF, segmentation, or optical flow to enhance control over head movement, gaze, and blinking (Xu et al., 2024a; Xiong et al., 2024; Zhou et al., 2021; Yang et al., 2024; Chatziagapi et al., 2024). For instance, VASA-1 (Xu et al., 2024a) utilizes 3D-aided representations in facial latent space to decouple features and generate high-quality faces. ConsistentAvatar (Yang et al., 2024) introduces a time-sensitive detail (TSD) map with a temporally diffusion module to align results with video frames. SegTalker (Xiong et al., 2024) decouples semantic regions into style codes, enabling talking segmentation driven by speech. PersonaTalk (Zhang et al., 2024a) uses cross-attention to inject speaking style into audio features, ensuring lip-sync accuracy. TalkinNeRF (Chatziagapi et al., 2024) integrates body posture, gestures, and facial expressions in a unified dynamic NeRF to generate animations with detailed hand and facial movements. However, these methods often suffer from limited intermediate representation capacity, which constrains video realism. Co-speech gesture generation further extends to the movement of gestures. ISCG (Ginosar et al., 2019) generates 2D skeletal gestures from audio and synthesizes them through a pose-to-image network. SDT (Qian et al., 2021) uses gesture template vectors combined with audio to create natural, synchronized upper body movements. ANGIE (Liu et al., 2022a) enhances 2D skeletal gestures by integrating learned template vectors and a gesture codebook, modeling body movements with unsupervised MRAA features, but often produces unnatural and inaccurate gestures. DiffTED (Hogue et al., 2024) improves temporal consistency and diversity in gestures using a thin-plate spline (TPS) motion model. Make-Your-Anchor (Huang et al., 2024) introduces shape constraints with 3D mesh conditioning and diffusion models, enabling precise torso and hand movements in anchor-style videos. Nevertheless, these methods still struggle with the generation quality of hand and lip regions and rely on multi-stage generation. In contrast, our method enables end-to-end generation without relying on other representations and focuses on regional supervised learning, enhancing the quality of specified parts.
2.2. Diffusion-based Video Generation
The diffusion-based model VDM (Ho et al., 2022) extends U-Net (Ronneberger et al., 2015) to capture temporal information for video generation. Make-A-Video (Singer et al., 2022) leverages pretrained text-to-image (T2I) models for efficient training without requiring paired text-video data. AnimateDiff (Guo et al., 2023) integrates a motion module for seamless use with T2I models, enabling temporally coherent animations. Beyond text-based methods, additional conditions like pose, skeletons, and audio enhance control. Animate Anyone (Hu, 2024) and DisCo (Wang et al., 2023a) use pose conditioners for motion guidance, while Champ (Zhu et al., 2025) employs SMPL for unified body representation with 2D skeleton guidance. MM-Diffusion (Ruan et al., 2023) achieves joint audio-video generation using dual U-Nets for aligned outputs. Our method operates in latent space, ensuring audio-video alignment with reduced inference costs.
3. Preliminary
Video Diffusion Models (VDM) (Chen et al., 2024; Guo et al., 2023; Wang et al., 2023b) extend image diffusion models for video generation by learning video distributions via denoising samples from a Gaussian distribution. A learnable autoencoder compresses videos into latent representations $z=E(x)$, and the diffusion model $\epsilon_{\theta}$ predicts noise $\epsilon$ at time $t$, conditioned on text $c_{text}$. The training objective is:
(1) $L_{video}=\mathbb{E}_{z,c,\epsilon\sim\mathcal{N}(0,I),t}\|\epsilon-\epsilon_{\theta}(z_{t},c_{text},t)\|^{2},$
Here, $z\in R^{F\times H\times W\times C}$ denotes the latent video code. $z_{t}$ is derived as $z_{t}=\lambda_{t}z_{0}+\sigma_{t}\epsilon$, with $\sigma_{t}=\sqrt{1-\lambda_{t}^{2}}$ controlling the diffusion scheduler. Studies (Peng et al., 2024) add control signals like image $c_{img}$ or audio $c_{a}$. During inference, the model denoises $z_{T}\sim\mathcal{N}(0,I)$ to $z_{0}$, and the frozen decoder reconstructs the video.
4. Method
This section introduces Parts-Aware Audio-Driven Human Animation (PAHA), an end-to-end audio-driven framework for generating half-body human animations with a diffusion model. Fig. 2 illustrates the overview of our PAHA. Given a speech audio $a$ and a reference character image $I_{\mathrm{ref}}$, the framework generates a motion video aligned with the audio.
The process is formulated as: $V=G_{\mathrm{PAR}}(I_{\mathrm{ref}},a,\nabla_{\mathrm{PCE}}(z_{t},a))$, where $G_{\mathrm{PAR}}(\cdot)$ denotes the unified video diffusion model trained with the PAR method, which simultaneously handles reference images and noisy videos without reference network. PAR identifies “Awareness Areas” we defined in key regions of the character (e.g., hands, face, body) in the training video frames based on a pose confidence-aware score, then apply dynamic loss weights in these areas to improve temporal smoothness and generation quality. $\nabla_{\mathrm{PCE}}(z_{t},a)$ represents the alignment guidance gradient produced during inference by diffusion-based regional classifiers trained using the PCE method. This gradient, based on noised latent features and audio conditions, ensures gestures and facial motions are consistent with the audio. The core process of PCE is the construction of classifiers and preparation of training samples. For inference, we design two guidance methods: Sequential Guidance for efficiency and Differential Guidance for quality. The upcoming sections cover the Unified Diffusion Model (Section 4.1.1), PAR (Section 4.1.2), PCE (Section 4.1.3), and the Inference Method (Section 4.2). Section 4.3 introduces the Chinese News Anchor Speech Dataset (CNAS), a co-speech Chinese dataset we created.
Figure 2. Overview of the proposed PAHA that consists of three core components: (a) The backbone of the Unified Video Diffusion Model (UniVDM) is a 3D U-Net. Video frames are encoded using a VAE encoder, while latent features of the reference image are extracted with both the CLIP and VAE encoders. These features are concatenated with the noisy input along the channel dimension, derived from either a conditioned first frame video or a noise video. (b) Parts-Aware Re-weighting (PAR) generates a dynamic loss re-weighting mask from confidence scores of video pose keypoints, improving supervision in specific regions during training. (c) Parts Consistency Enhancement (PCE) operates during inference, generating consistency gradients from audio and noise video features to enhance temporal visual-audio consistency.
4.1. Parts-Aware Audio-Driven Animation
4.1.1. Unified Video Diffusion Model (UniVDM)
We construct the diffusion model backbone network $G$ for video generation, as shown in Fig.2. To ensure temporally consistent character animation, we use the widely adopted 3D-UNet structure(Blattmann et al., 2023; Wang et al., 2024b). $G$ would denoise multi-frame noisy latent inputs into continuous video frames at each timestep, conditioned on a reference character image and driving audio.
Backbone Network. Diffusion-based video generation frameworks often use ControlNet-like 3D-UNet models (Zhang et al., 2023; Hu, 2024; Guo et al., 2023) to maintain temporal coherence, incorporating a reference encoder that replicates the 3D-UNet without temporal transformer layers to preserve the reference image’s details. However, these methods typically rely on multiple large networks, increasing parameter counts and optimization challenges.
To address this, we propose the unified video diffusion model (UniVDM), which processes reference images and noisy videos simultaneously by embedding reference information and estimating video content within a shared feature space. This design aligns features and ensures temporally coherent generation without requiring an additional reference encoder, reducing model parameters. The reference image is first encoded into latent space using a VAE encoder, producing a feature representation $f_{ref}$ with dimensions $C\times h\times w$ (channels, width, height). Next, the reference representation $f_{ref}$ and video features $f_{v}$ are stacked along the temporal dimension to form a merged feature $f_{merge}\in R^{(t+1)\times C\times h\times w}$, where $t$ is the temporal length. Finally, these combined features are processed by the unified diffusion model.
Audio layer. Audio is the primary signal driving the diffusion model $G$ to generate character animations. We create audio representation embeddings $A(f)$ for each frame by concatenating features from various modules of the pre-trained wav2vec (Schneider et al., 2019). Since motions can be influenced by future or past audio segments, such as mouth openings or inhalations before speaking, we define the speech features for each frame as: $A(f)=\oplus\{A(f-m),...A(f),...A(f+m)\}$, where $m$ represents the number of additional features on each side. To incorporate speech features into the generation process, we add audio attention layers after each reference attention layer in the backbone network, performing cross-attention between latent features $z_{t}$ and $A$: $z_{t}=\mathrm{CrossAttention}(z_{t},A(f))$.
4.1.2. Parts-Aware Re-weighting (PAR)
Experimental results show that while UniVDM effectively generates audio-driven half-body character animations, the quality of hand and face regions remains suboptimal. This suggests that relying solely on the pre-trained diffusion video model and cross-attention modules, $G$ faces challenges in fine-grained supervised learning for specific areas.
Motivation. The loss function of the diffusion model (Eq.1) assumes a uniform spatial prior; however, video frames often exhibit imbalances across foreground regions. Sequential frames can create uncertainties in dynamic appearance and motion, which negatively impact pose estimation and affect both training and inference. Furthermore, noisy pose guidance may lead to overfitting on misaligned samples, causing training instability.
To tackle these issues, we propose the Parts-Aware Re-weighting (PAR) method ( shown in Fig. 3), which allows for dynamic loss re-weighting of specific regions to improve foreground generation.
Awareness Area. Following the approach of ConvoFusion (Mughal et al., 2024), we divide the half-body character into three regions: hand, face, and body. Specifically, PAR leverages the confidence scores associated with each keypoint in the pose estimation model, where higher scores reflect better visual quality (less blur and occlusion). We establish a confidence score threshold $\tau_{j}$, considering keypoint $p_{i}$ with score $c_{p_{i}}$ above it as reliable. Then we generate loss re-weighting masks based on these thresholds. For pixels $x$ in smaller, dynamic regions (hand and face), circles with radius $r$ are drawn around reliable keypoints and merged for precise delineation. For the stable larger region (body), a rectangle is formed using the extreme x and y coordinates of all body keypoints. The “Awareness Area” $S_{\mathrm{a}}$ is thus defined as:
$S_{\mathrm{a}}&=\left\{\begin{array}[]{ll}\underset{c_{p_{i}}>\tau_{j}}{\cup}A(\text{circle}(p_{i},r)),&\text{if }p_{i}\in S_{\text{hand}}\cup S_{\text{face}}\\ \underset{c_{p_{i}}>\tau_{j}}{\cup}A(\text{rect}(x_{\text{min,max}},y_{\text{min,max}})),&\text{if }p_{i}\in S_{\text{body}}.\\ \end{array}\right.\end{aligned}$
Here, $A(\cdot)$ represents the area of the corresponding region, $\text{circle}(p_{i},r)$ denotes a circle with center $p_{i}$ and radius $r$. $\text{rect}(x_{\text{min,max}},y_{\text{min,max}})$ indicates a rectangle defined by the left and right x-coordinates $x_{\text{min,max}}$ and the top and bottom y-coordinates $y_{\text{min,max}}$.
Loss Re-weighting. During the video diffusion model’s loss computation, the “Awareness Area” $S_{\mathrm{a}}$ is assigned higher weights to prioritize its influence during training. For pixels $x$ in hand and face Awareness Areas, the mask weight is calculated by applying Gaussian smoothing to the sum of confidence scores $c_{p_{i}}$ of keypoints at circle centers, multiplied by the hyperparameter $\omega_{1}$. For body and other pixels $x$, the mask weight is set to the hyperparameter $\omega_{2}$, which defaults to 1 outside the body Awareness Area. The loss re-weighting mask $M(x)$ can be expressed as:
$M(x)&=\begin{cases}\mathcal{T}\left(\sum_{p_{i}\in S_{\text{hand}}\cup S_{\text{face}}}c_{p_{i}}\right)\cdot\omega_{1},&\text{if }x\in S_{\text{a}}^{\text{hand}}\cup S_{\text{a}}^{\text{face}}\\[4.30554pt] \omega_{2},&\text{if }x\in S_{\text{a}}^{\text{body}}\;\text{or else}.\end{cases}\end{aligned}$
Here, $\mathcal{T}(\cdot)$ represents the Gaussian smoothing operation. Following  (Wang et al., 2024a), which highlights poorer generation quality for small objects, $\omega_{2}$ is set lower than $\omega_{1}$ to better handle smaller areas. In summary, the PAR method focuses model training on specific regions, particularly the hands and face, enhancing visual quality and realism in generated content.
Figure 3. Process of our Parts-Aware Re-weighting (PAR) method. We identify the “Awareness Area” for the hand, face, and body regions based on pose keypoints and their confidence scores, then independently calculate weighted confidence scores for each region and merge them into a loss re-weighted mask.
4.1.3. Parts Consistency Enhancement (PCE)
Ensuring character motion aligns with driven audio is crucial for assessing framework performance. Most existing audio-visual strategies struggle to synchronize discrete character motion with the continuous audio spectrum, causing inherent video inconsistencies.
Motivation. A speaker’s body motions naturally coordinate with their spoken content, exhibiting rhythmic temporal correlations. However, we find that relying solely on single frames and long-distance audio for alignment can overlook temporal consistencies in visual features.
Thus, We propose the Parts Consistency Enhancement (PCE) method, which trains diffusion-based regional classifiers to discriminate between real and generated videos, implicitly learning local audio-video consistency. During inference, the classifier guides the process with gradients to synchronize key movements with speech.
Figure 4. Structure of our diffusion-based classifier. The pretrained diffusion encoder supports inputting noised video features and clean audio. After dimensionality reduction by GMP, the audio-video sequence is fed into the Transformer for full self-attention interaction, and the MLP head finally predicts the audio-video synchronization score.
Classifier Construction. After training UniVDM, We start to train audio-visual classifiers combining a diffusion model U-Net encoder and a transformer encoder, as shown in Fig. 4. The input consists of noised video latent features $V_{i}$ and clean audio samples $A_{j}$. Prior works (Xiao et al., 2022; Sauer et al., 2025; Lin et al., 2024) have used adversarial discriminators in diffusion model distillation to accelerate sampling. We choose the U-Net encoder from the pre-trained Unified Diffusion Model (Section 4.1.1) as the classifier’s initial module, offering two advantages: 1) leveraging the diffusion model’s understanding of audio-video modalities to simplify training; 2) processing latent features across all diffusion timesteps, avoiding pixel-space mapping and reducing computational costs.
Building on the above discovery regarding frequency domain and motion relationships, we need to extract time-domain features. Drawing inspiration from natural language translation tasks, we consider the correlation between motions and spectral information to be analogous to the relationship between “vocabulary” and “sentence”. Latent video features $f_{v}\in\mathbb{R}^{c\times h\times w\times t_{v}}$ are concatenated with clean audio features $f_{a}\in\mathbb{R}^{c\times t_{a}}$ and serialized them into the transformer (Vaswani, 2017) encoder for interaction through self-attention.
Directly flattening all visual features and densely combining them with audio features for the transformer is computationally expensive, with a quadratic complexity of $\mathcal{O}((hwt_{v}+t_{a})^{2})$, limiting scalability for longer videos. To address this, we apply global max pooling (GMP) to each video frame instead of dense visual inputs. We also introduce a learnable class token ([CLS]) to help the classifier distinguish modalities while preserving spatial-temporal positional information. We add modal encoding $E_{m}\in\mathbb{R}^{c\times 2}$ to the video and audio features, indicating the feature type (i.e., audio or visual), along with temporal encoding $E_{t_{\{v,a\}}}\in\mathbb{R}^{c\times t_{\{v,a\}}}$. The video features also include 3D RoPE (Su et al., 2024) positional encoding embeddings $E_{s}\in\mathbb{R}^{c\times h\times w}$, which are efficient for varying video token counts. This process is expressed as:
(4) $\overline{V}_{i}$ $=\text{GMP}(V_{i})+E_{m}+E_{t_{v}}+E_{s},$
(5) $\overline{A}_{j}$ $=A_{j}+E_{m}+E_{t_{a}},$
[…]
Here, $\oplus$ indicates a concatenation operation. The length of the input sequence $Z_{ij}$ to the transformer encoder is reduced from $(hwt_{v}+t_{a}+1)$ to $(t_{v}+t_{a}+1)$, resulting in significant memory savings. A multi-layer perception (Taud and Mas, 2018) serves as the classifier $h$.
Ensuring character motion aligns with driven audio is as crucial as video quality for evaluating framework performance. Many existing audio-visual methods fail to synchronize discrete character motions with continuous audio, causing video inconsistencies.
Training Data. Inspired by Diffusion Self-Distillation (Cai et al., 2024), we generate negative samples using the pre-trained unified video diffusion model $G$, while positive samples come from real videos, as shown in Fig. 5. We employ a masking strategy to help the classifier focus on consistent features between motion video and co-speech audio. We train two classifiers: non-face classifier and face classifier. To train the non-face(i.e., areas other than the face) classifier, facial regions are randomly masked to minimize their influence on audio matching, emphasizing non-face motions. For the face classifier, only facial regions are retained while masking the rest. Additionally, we use two simple data augmentation strategies: 1) random reference frame sampling during inference, starting with the first frame, and 2) varying audio and video lengths. These strategies improve the one-to-many mapping between audio and motion.
Training Loss. We use only the first token output from the final encoder layer ($Y^{1}_{ij}$), corresponding to the [CLS] position, as the aggregated representation of the entire output sequence for the MLP. The output is a synchronisation score $s_{ij}$, indicating to what degree the inputs $V_{i}$ and $A_{i}$ are in sync, $s_{ij}=h(Y^{1}_{ij})$. We optimize the classifier using binary cross-entropy loss:
(7) $L_{cls}=-(y\log(s)+(1-y)\log(1-s)).$
Here, $y$ represents the final predicted label.
4.2. Inference Process
The face and non-face classifier are trained separately offline but could work simultaneously during inference, each focusing on its assigned area. We propose two effective inference methods.
4.2.1. Sequential Guidance (SG)
Start by sampling $z_{T}\sim\mathcal{N}(0,I)$ from a Gaussian distribution. During inference, the numerical solver $f_{\theta}$ first predicts $\hat{z}_{t}=f_{\theta}(\hat{z}_{t+1},a,t+1)$. Then the non-face classifier $D_{non\text{-}face}$ computes the gradient $\nabla D_{non\text{-}face}(\hat{z}_{t},a)\in\mathbb{R}^{c\times h\times w}$ conditioned on $(\hat{z}_{t},a)$. This gradient spatially influences only the potential features guided by the non-face mask area $M_{non\text{-}face}$, resulting in:
$\hat{z}_{t}^{non\text{-}face}=\hat{z}_{t}-\lambda_{non\text{-}face}\sigma_{t}\nabla D_{non\text{-}face}(\hat{z}_{t},a)*M_{non\text{-}face},$
where $\lambda_{non\text{-}face}$ is the gradient weight used to control the strength of the condition. Next, $(\hat{z}_{t}^{non\text{-}face},a)$ is used as the condition for the facial classifier $D_{face}$ to compute the facial consistency gradient $\nabla D_{face}(\hat{z}^{non\text{-}face}_{t},a)\in\mathbb{R}^{c\times h\times w}$. Similarly, by introducing $M_{face}$, this gradient works only on the facial region:
(9) $\hat{z}_{t}^{face}=\hat{z}_{t}-\lambda_{face}\sigma_{t}\nabla D_{face}(\hat{z}^{non\text{-}face}_{t},a)*M_{face}.$
The video latent $\hat{z}_{t}^{*}$ is derived through classifier-based guidance, yielding the final predicted clean sample $\hat{z}_{0}$ via above iterative process.
(10) $\hat{z}_{t}^{*}$ $=\hat{z}_{t}^{face}+\hat{z}_{t}^{non\text{-}face}-\hat{z}_{t},$
(11) $\hat{z}_{t-1}=f_{\theta}(\hat{z}^{*}_{t},a,t).$
Figure 5. Pipeline for constructing negative samples for the classifier. The video is generated by our pre-trained UniVDM, conditioned on randomly sampled audio and reference frames from the real dataset. We mask specific areas with a certain probability to enhance local modality alignment.
4.2.2. Differential Guidance (DG)
In our experiments, we observe that classifiers negatively affect the consistency of non-guided areas; for instance, the facial classifier reduces non-face alignment metrics (Table. 3 ), and even mask conditions fails to completely isolate this effect. This is caused by Eq.11, where gradients from face and non-face areas interact through the diffusion model.
To resolve this, we propose differential guidance to counteract negative changes caused by classifiers in non-guided areas. Specifically, we revise Eq.11 in Sequential Guidance to:
$\hat{z}_{t-1}=$ $f_{\theta}(\hat{z}_{t}^{*},a,t)+\lambda_{diff}(f_{\theta}(\hat{z}_{t}^{non\text{-}face},a,t)*M_{face}$
(12) $+f_{\theta}(\hat{z}_{t}^{face},a,t)*M_{non\text{-}face}).$
Here, $\lambda_{diff}$ is the weight that controls the differential strength. The formula can be understood as a compensation for the non-guided areas of the corresponding classifier. Differential Guidance, while slower than Sequential Guidance during inference, achieves superior generation quality (Table. 1). The pseudocode for SG, DG, and the inference pipeline is detailed in Appendix B.4.2.
4.3. CNAS Dataset
To address the scarcity of Chinese co-speech data, we construct the Chinese News Anchor Speech Dataset (CNAS). The dataset features single speakers primarily facing the camera, with above-waist perspectives and communication in Chinese. After preprocessing, 1,473 valid clips are obtained. More details are provided in Appendix C. This Chinese broadcasting dataset will be made publicly available for broader research.
5. Experiments
Table 1. Quantitative results on the pats and cnas datasets. the best metrics are highlighted in BLOD and UNDERLINE indicates the second-best. SG stands for Sequential Guidance, and DG stands for Differential Guidance.
Method PATS CNAS
FVD $\downarrow$ BAS $\uparrow$ FGD $\downarrow$ Sync-C $\uparrow$ Div. $\uparrow$ FVD $\downarrow$ BAS $\uparrow$ FGD $\downarrow$ Sync-C $\uparrow$ Div. $\uparrow$
SDT (Qian et al., 2021) 2281.65 0.1067 22.966 5.337 83.917 1592.48 0.0819 28.559 3.374 29.73
ANGIE (Liu et al., 2022a) 2477.23 0.0892 42.081 4.927 79.102 / / / / /
MM-Diffusion (Ruan et al., 2023) 2204.09 0.1085 23.782 5.348 82.545 1531.21 0.0883 27.664 3.288 26.11
S2G (He et al., 2024) 2118.41 0.1227 19.816 5.294 85.623 1465.75 0.1087 27.523 3.452 32.65
PAHA-SG (Ours) 2052.95 0.1561 15.021 6.183 100.917 1379.57 0.1241 24.439 3.81 39.18
PAHA-DG (Ours) 2048.75 0.1602 14.661 6.328 98.235 1362.44 0.1275 23.715 4.092 37.54
Table 2. Ablation study regarding guidance time. in each number pair, the former is PAHA-SG and the latter is PAHA-DG. indicates the best of PAHA-SG, indicates the best of PAHA-DG.
Guidance Time Rate FVD $\downarrow$ BAS $\uparrow$ FGD $\downarrow$ Sync-C $\uparrow$ Div. $\uparrow$ TC $\downarrow$
0% 2047.53 / 2047.53 0.1346 / 0.1346 17.675 / 17.675 5.709 / 5.709 90.804 / 90.804 85 / 85
25% 2050.66 / 2048.09 0.1423 / 0.1507 16.790 / 16.238 6.015 / 6.094 93.713 / 95.482 91 / 96
50% 2052.95 / 2048.75 0.1561 / 0.1602 15.021 / 14.661 6.183 / 6.328 99.604 / 98.235 98 / 112
75% 2056.03 / 2051.78 0.1662 / 0.1689 17.882 / 17.104 6.306 / 6.451 96.122 / 97.318 107 / 129
100% 2062.95 / 2055.43 0.1674 / 0.1711 18.723 / 17.868 6.382 / 6.546 92.071 / 93.944 114 / 141
5.1. Experimental Settings
Implement details. UniVDM is initialized from a pretrained video diffusion model (Blattmann et al., 2023). During training with the PAR method, videos have a spatial resolution of 512×512 and a fixed length of 32 frames. We choose DDPM (Ho et al., 2020) as the noise scheduler with 1000 sampling steps. The confidence score threshold $\tau_{j}$ is set to 0.8, with a radius $r$ of 10. The hand/face weight hyperparameter $\omega_{1}$ is set to 10, while $\omega_{2}$ for the body region is set to 2. In PCE classifiers training, negative samples are generated with a 30-step DDIM (Song et al., 2020) sampler. Facial boxes are extracted using MediaPipe (Lugaresi et al., 2019). Both the non-face classifier and face classifier encoders are initialized with pre-trained $G$. For final inference, we use a 30-step DDIM sampler, applying classifier guidance only in the first 15 steps. More implementation details can be found in Appendix B.
Datasets. The training data for our backbone network $G$ comes from the PATS dataset (Ahuja et al., 2020) and the CNAS dataset we propose. For fair comparison, we use the same training subset as S2G (He et al., 2024), which includes four talkers: Jon, Chemistry, Oliver, and Seth. Each speaker has 1,200 valid clips (without cutouts or camera movements), with clip lengths ranging from 4 to 15 seconds, at 25 fps, and training resolution uniformly adjusted to 512$\times$512. 90% of the data is used for training and 10% for evaluation. The CNAS dataset follows the same configuration as PATS. Positive audio-video samples for training the PCE classifiers are sourced from the processed PATS training set, while negative samples are generated by the backbone network $G$ pretrained for 60k steps, creating one-to-one paired positive and negative samples. Data augmentation strategies from Section 4.1.3 expand the training set, with audio and video lengths uniformly sampled across 30, 60, 90, and 120 frames. These augmentations yield 50k audio-video pairs, with a mask probability of 80%.
Evaluation metrics. 1) Fréchet Gesture Distance (FGD) (Yoon et al., 2020); 2) Diversity (Div.); 3) Beat Alignment Score (BAS) (Li et al., 2021) 4) Synchronization-C (Sync-C) (Xu et al., 2024b); 5) Fréchet Video Distance (FVD) (Unterthiner et al., 2018). The meanings of each metric are detailed in Appendix B.5.
Baselines. We compare our method with four baselines: 1) S2G (He et al., 2024), the latest SOTA in gesture video generation; 2) ANGIE (Liu et al., 2022a); 3) SDT (Qian et al., 2021); and 4) MM-Diffusion (Ruan et al., 2023). All baselines are initialized with official weights and fine-tuned on the PATS and CNAS datasets.
5.2. Quantitative Results
The quantitative results reported in Table 1 show that our method achieves the best performance in quality, alignment, and motion metrics. This demonstrates that our end-to-end Unified Diffusion Model can generate realistic motion videos with high speech consistency under the simultaneous influence of the PAR and PCE methods. Our approach using Differential Guidance outperforms other metrics except Div. compared to Sequential Guidance. Furthermore, even though our method primarily focuses on optimizing specific areas, the improvement in FVD strongly demonstrates the method’s gain in overall quality.
5.3. User Study
Audio-driven human animation relies heavily on subjective perception over objective metrics. To evaluate our method’s visual performance, we conducted a user study comparing videos generated by S2G, MM-Diffusion, SDT, and PAHA-DG. Ten videos were randomly selected from the PATS and CNAS test sets, and 16 participants evaluated them based on realism, diversity, speech-motion synchronization, and overall quality. Participants were instructed to ignore texture and facial expressions during motion evaluations. Preferences for each criterion were measured independently. As shown in Table 4, our method outperformed others across all criteria, especially in overall quality and synchronization, proving its ability to generate high-quality co-speech gesture videos while balancing motion and visual effects.
Table 3. Ablation study regarding core modules. BOLD indicates the best “w/o” is short for “without”.
Name FVD $\downarrow$ BAS $\uparrow$ FGD $\downarrow$ Sync-C $\uparrow$ Div. $\uparrow$
w/o PAR 2145.72 0.1497 15.452 5.977 89.866
w/o non-face classifier 2087.49 0.1302 17.957 6.583 97.483
w/o face classifier 2060.11 0.1554 15.181 5.726 99.154
w/o PAR+two classifiers 2126.72 0.1197 20.052 5.227 83.772
w/o augmentation 2068.74 0.1548 15.279 6.126 99.604
Ours 2052.95 0.1561 15.021 6.183 100.917
5.4. Ablation Study
Core Components. We identify three key components in the method: PAR, non-face classifier guidance, and face classifier guidance. Sequential Guidance is used for video generation during inference. Table 3 shows that removing PAR worsens video quality (FVD +92.77) and reduces Diversity (-11.051). Face classifier guidance significantly affects lip alignment (Sync-C -14.9%), while the non-face classifier primarily impacts gesture movement (FGD +19.5%, BAS -16.6%), aligning with expectations. Both classifiers improve alignment and motion metrics but slightly degrade FVD, with the face classifier having a stronger effect.
Guidance Parameters. We determine the guidance strength during inference via ablation experiments. The complete results for the non-face guidance weight $\lambda_{non\text{-}face}$, face guidance weight $\lambda_{face}$ and differential guidance weight $\lambda_{diff}$ are available in Appendix D. The final weight combination is set to $(\lambda_{non\text{-}face}=1,\lambda_{face}=0.1,\lambda_{diff}=0.25)$, achieving optimal performance.
Guidance Time Rate. Table 2 compares the performance of PAHA-SG (Sequential Guidance) and PAHA-DG (Differential Guidance) at different proportions of guidance steps. A new metric, Time Cost (TC), measures the average time (in seconds) to generate a video segment (256x256 resolution, 30 steps), rounded to the nearest integer. In the paired values, the first represents PAHA-SG, and the second is PAHA-DG. The data shows that increasing classifier-guided steps raises time costs. While FVD worsens, other metrics improve. The best performance balance occurs as the guided step proportion increases from 0% to 50%, but further increases reduce performance.
Inference Time. As shown in Fig. 1, our method significantly outperforms the multi-stage co-speech gesture method in inference efficiency using the TC metric. While its inference time is comparable to ANGIE, our method delivers noticeably better results. Operating in the latent space further enhances its efficiency.
Data Augmentation. We also perform an ablation analysis of the data augmentation strategy, which demonstrates its effectiveness in improving all metrics (Table. 3).
Table 4. The preferred percentage of our method and the baselines in user study on the PATS and CNAS dataset.
Method Realness $\uparrow$ Diversity $\uparrow$ Synchrony $\uparrow$ Overall quality $\uparrow$
SDT (Qian et al., 2021) 15.00% 8.20% 10.00% 6.20%
MM-Diff (Ruan et al., 2023) 8.60% 13.60% 11.80% 10.60%
S2G (He et al., 2024) 12.60% 15.60% 13.60% 13.80%
PAHA-DG 63.80% 62.60% 64.60% 69.40%
6. Conclusion
This paper introduces PAHA, an end-to-end diffusion-based framework for generating high-quality audio-driven half-body human animations without intermediate representations. The unified video diffusion model processes reference images and noisy videos simultaneously. During training, Parts-Aware Re-weighting (PAR) dynamically adjusts loss to focus on foreground regions, enhancing visual quality. The Parts Consistency Enhancement (PCE) method uses self-distillation to train diffusion-based regional classifiers, which provide alignment gradients during inference to ensure motion-audio consistency in specific areas. Experiments show that PAHA produces visually appealing, temporally consistent animations, surpassing existing methods. Additionally, we present CNAS, the first Chinese broadcasting dataset for co-speech gestures.
Appendix
Appendix A Overview
In this supplementary material, more details about the proposed PAHA and more experimental results are provided, including:
Appendix B Additional Implementation Details of PAHA
B.1. UniVDM
UniVDM is initialized from a pretrained video diffusion model (Blattmann et al., 2023). During training with the PAR method, videos have a spatial resolution of 512×512, a fixed length of 32 frames, a batch size of 8, and 100k training steps. We use the AdamW optimizer (Loshchilov, 2017) with a learning rate of 5e-5 and DDPM (Ho et al., 2020) as the noise scheduler with 1000 sampling steps. The confidence score threshold $\tau_{j}$ is set to 0.8, with a radius $r$ of 10. The hand/face weight hyperparameter $\omega_{1}$ is set to 10, while $\omega_{2}$ for the body region is set to 2.
B.2. PAR
B.2.1. Training Details
When training the backbone network $G$ using the PAR method, we use DWpose (Yang et al., 2023) to extract the required pose sequences. The visual encoder from the multimodal CLIP-Huge model (Sohl-Dickstein et al., 2015) in Stable Diffusion v2.1 is used to encode CLIP embeddings of reference images.
B.3. PCE
B.3.1. Motivation
Fig. 6 shows the correlation between motions and the spectrogram between spectral energy changes and localized motions (hands, lips). When the speaker begins to talk, the mid-to-high frequencies in the spectrum brighten, while darker areas of the spectrum correspond to smaller motion amplitudes.
Figure 6. The correlation between motions and the corresponding spectrogram. When the speaker begins to talk, the mid-to-high frequencies in the spectrum brighten, while darker areas of the spectrum correspond to smaller motion amplitudes.
B.3.2. Dataset Construction for Classifier
We trained the PCE using offline training, with positive samples coming from the ground truth and negative samples generated using the pre-trained UniVDM (60k steps checkpoints).
B.3.3. Training Details
In the PCE classifier training experiments, all negative samples are generated using a DDIM (Song et al., 2020) sampler, with 30 sampling steps. We use MediaPipe (Lugaresi et al., 2019) to obtain facial boxes. Both the non-face classifier $D_{non\text{-}face}$ and face classifier $D_{face}$ use a learning rate of 1e-4, with a batch size of 16 and 100k training steps. A warmup strategy is applied for the first 5000 steps, and the encoders of both classifiers are initialized with pre-trained $G$ checkpoints. When training the classifier, we used the same noise scheduler (DDPM) as when training the Unified Video Diffusion Model (UniVDM).
B.4. Inference
B.4.1. Long Video Generation
Memory constraints make generating long videos in a single pass unfeasible. Thus, multiple short video segments must be synthesized separately and merged. Existing methods typically use a sliding window strategy to synthesize short videos from overlapping local windows and merge them by averaging the overlaps, but this can cause segment discontinuities.
In our Unified Diffusion Model, we propose a unified noise input that allows random noise videos or first-frame conditioned videos as input for video synthesis. The first-frame conditioning method uses the initial frame as the condition for generating videos starting from the frame. By utilizing this strategy, the last frame of the previous short video segment can serve as the first frame of the next segment, achieving seamless and visually coherent long animation.
B.4.2. Algorithm
For final inference, we use a 30-step DDIM sampler, applying classifier guidance only in the first 15 steps. The guidance strengths for the face classifier, non-face classifier, and differential guidance ($\lambda_{face}$, $\lambda_{non\text{-}face}$, $\lambda_{diff}$) are set to 0.1, 1, and 0.25, respectively. We set $\lambda_{face}$ significantly lower than $\lambda_{non\text{-}ace}$ because experiments show that the face classifier tends to dominate the generation process more easily than the non-face classifier, and excessive $\lambda_{face}$ notably decreases video quality. Furthermore, we observe that both classifiers improved alignment and motion metrics but also caused varying degrees of FVD decline, with the face classifier having a more pronounced impact. Therefore, in Sequential Guidance (SG), we first use the non-face classifier and apply a smaller guidance weight to the face classifier to mitigate its influence. The evaluation resolution of the final video is 256$\times$256.
Figure 7 illustrate the inference pipelines for PAHA-SG(Sequential Guidance) and PAHA-DG(Differential Guidance). The only difference between the two inference methods lies in the final calculation of $\hat{z}_{t_{n-1}}$. In SG, gradients from the face and non-face classifiers are sequentially computed and fed back into the denoising video. In contrast, DG performs two additional ODE Solver $f_{\theta}(\cdot)$ executions to mitigate classifier negative impact in non-guided regions. The pseudo code of our classifier-based inference process is shown in Algorithm 1.
Figure 7. The inference pipeline of PAHA includes two forms: PAHA-SG (Sequential Guidance) and PAHA-DG (Differential Guidance), focusing on generation efficiency and quality, respectively.
Algorithm 1 Classifier-based Inference
0:  A reference image $\mathbf{I_{ref}}$, a driven-audio $\mathbf{a}$, the unified diffusion model $G_{\mathrm{PAR}}$ trained by PAR, the face classifier $D_{face}$, the non-face classifier $D_{non\text{-}face}$, timestep $t_{N-1}>t_{N-2}>...>t_{1}$, ODE solver $f_{\theta}$, noise scheduler $\alpha(t),\sigma(t)$, face mask $M_{face}$, non-face mask $M_{non\text{-}face}$, guidance weight $\lambda_{face},\lambda_{non\text{-}face},\lambda_{diff}$, guidance rate $\mathbf{r}$, decoder $\mathbf{D}$
0:  the generated co-speech video $\mathbf{v}$
1:  Sample Guassian Noise $z_{T}\sim\mathcal{N}(0,I)$
2:  for $n=N-1$ to $(N-1)(1-r)$ do
3:     $\hat{z}_{t_{n}}=f_{\theta}(\hat{z}_{t_{n+1}},a,t_{n+1})$
4:     $\hat{z}_{t_{n}}^{non\text{-}face}=&\hat{z}_{t_{n}}-\lambda_{non\text{-}face}\sigma_{t_{n}}\nabla D_{non\text{-}face}(\hat{z}_{t_{n}},a)*\\ &M_{non\text{-}face}\end{aligned}$
5:     $\hat{z}_{t_{n}}^{face}=\hat{z}_{t_{n}}-\lambda_{face}\sigma_{t_{n}}\nabla D_{face}(\hat{z}^{non\text{-}face}_{t_{n}},a)*M_{face}$
6:     $\hat{z}_{t_{n}}^{*}=\hat{z}_{t_{n}}^{face}+\hat{z}_{t_{n}}^{non\text{-}face}-\hat{z}_{t_{n}}$
7:     if guidance name is “Sequential Guidance (SG)” then
8:        $\hat{z}_{t_{n-1}}=f_{\theta}(\hat{z}^{*}_{t_{n}},a,t_{n})$
9:     else if guidance name is “Differential Guidance (DG)” then
10:        $\hat{z}_{t_{n-1}}=&f_{\theta}(\hat{z}_{t_{n}}^{*},a,t_{n})+\lambda_{diff}(f_{\theta}(\hat{z}_{t_{n}}^{non\text{-}face},a,t_{n})*M_{face}+\\ &f_{\theta}(\hat{z}_{t_{n}}^{face},a,t_{n})*M_{non\text{-}face})\end{aligned}$
11:     end if
12:  end for
B.5. Evaluation Metrics
To evaluate the quality, diversity, and alignment between gestures and speech, we employ: 1) Fréchet Gesture Distance (FGD) (Yoon et al., 2020), measuring the distribution gap between real and generated gestures in feature space; 2) Diversity (Div.), calculating the average feature distance between generated gestures. These metrics use an autoencoder trained on PATS and CNAS poses following the code from (Liu et al., 2022b). Additionally, in accordance with S2G, we calculate the 3) Beat Alignment Score (BAS) (Li et al., 2021), measuring the average distance between speech and gesture beats. 4) Synchronization-C (Sync-C) (Xu et al., 2024b) evaluates lip synchronization in generated videos, where a higher score reflects better alignment of lip motions with the audio. For video evaluation, we use 5) Fréchet Video Distance (FVD) (Unterthiner et al., 2018) to assess the overall quality of gesture videos, computed in feature space using an I3D classifier (Wang et al., 2019) pre-trained on Kinetics-400 (Kay et al., 2017).
Appendix C More Details about CNAS
With the goal of modeling human bodies, we estimate joints of 2D body and hands, and fit statistical 2D human body models by minimizing projection errors and temporal differences between consecutive frames. We further employ the advanced detection model DINO v2 (Oquab et al., 2023) to assist with body part detection. We filter out videos with significant screen switching, undetected or partially detected faces or bodies, unstable detection results, and poor audio quality. This process generates a dataset of 36,825 seconds with 5 identity IDs, containing 1,473 valid clips per news anchor, with a video resolution of 512$\times$896. Each clip contains 125 frames at 25 fps (5 seconds) with audio sampled at 22 kHz.
Appendix D More Ablation Study
Guidance Parameters. We determine the guidance strength during inference via ablation experiments. Tables 65, and 7 present the quantitative results for non-face guidance weight $\lambda_{non\text{-}face}$, face guidance weight $\lambda_{face}$, and differential guidance weight $\lambda_{diff}$ when executing Differential Guidance during the inference. The final weight combination is set to $(\lambda_{non\text{-}face}=1,\lambda_{face}=0.1,\lambda_{diff}=0.25)$, achieving optimal performance.
Table 5. Ablation study regarding non-face guidance strength. BOLD indicates the best.
$\lambda_{non\text{-}face}$ FVD $\downarrow$ BAS $\uparrow$ FGD $\downarrow$ Sync-C $\uparrow$ Div. $\uparrow$
0.8 2049.91 0.1578 14.107 6.436 95.271
0.9 2047.68 0.1591 14.393 6.390 96.648
1.0 2045.75 0.1602 14.661 6.328 98.235
1.1 2044.09 0.1599 14.917 6.247 101.443
1.2 2042.87 0.1596 15.265 6.146 100.143
Table 6. Ablation study regarding face guidance strength. BOLD indicates the best.
$\lambda_{face}$ FVD $\downarrow$ BAS $\uparrow$ FGD $\downarrow$ Sync-C $\uparrow$ Div. $\uparrow$
0.05 2041.67 0.1629 14.682 6.197 100.267
0.10 2045.75 0.1602 14.661 6.328 98.235
0.15 2047.85 0.1594 14.736 6.416 96.563
0.20 2050.99 0.1581 14.799 6.480 94.782
0.25 2056.39 0.1571 14.882 6.529 93.551
Table 7. Ablation study regarding differential guidance strength. BOLD indicates the best.
$\lambda_{diff}$ FVD $\downarrow$ BAS $\uparrow$ FGD $\downarrow$ Sync-C $\uparrow$ Div. $\uparrow$
0.15 2053.63 0.1584 14.805 6.257 99.580
0.20 2048.11 0.1596 14.739 6.301 98.742
0.25 2045.75 0.1602 14.661 6.328 98.235
0.30 2043.98 0.1593 14.582 6.296 97.611
0.35 2045.53 0.1582 14.355 6.235 97.078
Appendix E Limitations and Future Work
Although the proposed method PAHA improves regional generation quality and motion-audio synchronization, our categorization of optimized areas remains broad, encompassing only face, hands, and body lacks finer control. This results in the model still needing improvement in realistically rendering complex expressions (emotional changes) and complex gestures (intersections and overlaps). Meanwhile, the fixed perspective of generated videos indicates the model’s limitations in handling dynamic scenes. Furthermore, our Chinese co-speech dataset CNAS remains smaller than mainstream English datasets (PATS) in terms of character IDs and the diversity of actions and expressions. However, the experimental results show that although the dataset is small in scale, it is sufficient to validate the algorithm’s effectiveness.
In the future, we plan to develop real-time feedback mechanisms to enhance the interactivity and realism of human animation, improving the model’s robustness across viewpoints and interactions for broader applications in live media and augmented reality. Additionally, we aim to incorporate diverse speaking styles and emotions to enhance expressiveness and control, and further expand the Chinese co-speech dataset CNAS.